Calculate the mass percent of 2.57 g KCl

2.57 g KCl dissolved in 40.6 g H2O.



1
Expert's answer
2020-10-12T13:27:18-0400

W(KCl)=m(KCl)/(m(KCl)+m(H2O))=2.57/(2.57+40.6)=2.57/43.17=0,0595=5,95%

Answer: W(KCl)=5.95%